# Meisam Kabiri1, and Holger Voos12 1 Interdisciplinary Center for Security Reliability and Trust (SnT)

### v3

@@ IV-C Augmenting the Dataset with Simulated 5G ToA Measurements
Dataset Frequency (GHz) Base Station Std (cm) Mean (cm)
Dataset0 28 1 $32.92$ $4.20$
2 $36.05$ $-1.21$
3 $31.99$ $4.19$
4 $41.32$ $8.28$
78 1 $19.58$ $-2.71$
2 $19.44$ $1.34$
3 $19.10$ $0.31$
4 $14.25$ $-1.72$
Dataset1 28 1 $36.31$ $-6.88$
2 $29.60$ $-3.56$
3 $36.60$ $1.88$
4 $32.22$ $-2.68$
78 1 $15.95$ $-2.83$
2 $17.01$ $-2.29$
3 $16.08$ $0.56$
4 $15.64$ $-0.86$
Dataset2 28 1 $38.44$ $-8.24$
2 $31.94$ $-4.93$
3 $33.25$ $7.97$
4 $30.49$ $-0.50$
78 1 $15.83$ $0.87$
2 $18.92$ $-4.97$
3 $16.38$ $-0.80$
4 $16.63$ $-0.59$
Dataset3 28 1 $40.14$ $-14.56$
2 $32.19$ $-6.68$
3 $31.17$ $4.49$
4 $27.64$ $-4.97$
78 1 $15.03$ $0.87$
2 $17.61$ $-8.71$
3 $14.69$ $-1.49$
4 $15.94$ $-1.06$
Dataset4 28 1 $38.23$ $-19.41$
2 $29.54$ $-2.18$
3 $31.23$ $-3.09$
4 $40.62$ $-8.31$
78 1 $14.40$ $3.12$
2 $16.80$ $-4.67$
3 $15.56$ $0.57$
4 $18.76$ $-1.08$
That table reveals differences between the two frequency bands across all datasets. The 78 GHz configuration demonstrates superior precision with standard deviations ranging from 14.25-19.58 cm, compared to the 28 GHz measurements which show larger variations of 27.64-41.32 cm. Mean errors at 78 GHz typically remain within ±5 cm, while 28 GHz measurements exhibit larger systematic biases up to 19.41 cm. The 78 GHz configuration’s superior performance, evidenced by approximately 50% lower standard deviation and smaller systematic biases across all datasets, suggests it is the preferred choice for precise ToA measurements in indoor drone navigation scenarios. This performance advantage can be attributed to the larger bandwidth (400 MHz vs 200 MHz) and higher subcarrier spacing (240 kHz vs 120 kHz) available at 78 GHz.
IV-E Evaluation Metrics
We evaluated the system using the Absolute Trajectory Error (ATE) metric, comparing the estimated trajectories against the OptiTrack system’s ground truth. Two distinct evaluation approaches were employed:
  • Local Evaluation: Before computing the ATE, we estimate and apply an optimal transformation (rotation, translation, and scale) between the SLAM-estimated trajectory and the ground truth. This alignment process factors out global positioning errors and evaluates the system’s ability to maintain consistent local trajectory shape and motion patterns. For example, if the drone flies in a square pattern, this evaluation would measure how well the SLAM system captures the square’s shape, regardless of its absolute position or orientation in space.
  • Global Evaluation: No transformation or scale optimization was applied, enabling direct comparison between estimated and ground truth trajectories in their original coordinate frames. This approach provides a stringent assessment of the system’s ability to maintain accurate absolute positioning and orientation across the entire flight path.
Both evaluations were compared against the baseline ORB-SLAM3 system without ToA data integration, demonstrating how 5G measurements enhance global positioning capabilities while preserving local consistency.
@@ V Results
In this section, we present the experimental results that evaluate the performance of different SLAM modes integrated with 5G ToA data at two frequencies: 28 GHz and 78 GHz. Table II presents comprehensive performance metrics for four SLAM configurations: RGB-D, RGB-D Inertial, Monocular (MONO), and Monocular Inertial. The results are evaluated across five distinct datasets (Dataset0 through Dataset4) collected in the Aerolab environment. For each configuration and dataset, we report three key metrics: local error (in meters), global error (in meters), and local improvement percentage compared to the baseline. To ensure statistical reliability, each experiment was conducted five times, and the values presented in the table represent the average ATE across these runs. The table includes both baseline measurements (without ToA integration) and results with ToA integration at both frequency bands, along with mode averages for each SLAM configuration.
TABLE II: Average Errors and Improvements for Different SLAM Modes. Note: For the monocular mode with TOA integration, results for Datasets 2 and 4 are calculated up to 75 s and 65 s respectively, as the system lost tracking beyond these durations.
SLAM Mode Dataset Configuration Local (m) Global (m) Local Imp.
RGB-D Dataset0 Baseline 0.064 N/A
@@
RGB-D Inertial Dataset0 Baseline 0.089 N/A
@@
28 GHz ToA 0.094 0.280 0.42%
78 GHz ToA 0.090 0.262 5.97%
MONO Dataset0 Baseline Failed N/A
@@ 78 GHz ToA Failed Failed
Dataset1 Baseline 1.370 N/A
28 GHz ToA 0.059 0.380 95.7%
78 GHz ToA 0.058 0.121 95.8%
Dataset2 Baseline 1.284 N/A
28 GHz ToA 0.089 0.186 93.1%
78 GHz ToA 0.075 0.152 94.2%
Dataset3 Baseline 1.363 N/A
28 GHz ToA 0.058 0.162 95.7%
78 GHz ToA 0.058 0.127 95.7%
Dataset4 Baseline 1.041 N/A
28 GHz ToA 0.179 0.320 82.8%
78 GHz ToA 0.061 0.131 94.1%
Mode Average Baseline 1.265 N/A
28 GHz ToA 0.096 0.262 92.4%
78 GHz ToA 0.063 0.133 95.0%
MONO Inertial Dataset0 Baseline Failed N/A
28 GHz ToA 0.096 0.434 INF
78 GHz ToA 0.076 0.306 INF
Dataset1 Baseline Failed N/A
28 GHz ToA 0.112 0.190 INF
78 GHz ToA 0.121 0.370 INF
Dataset2 Baseline Failed N/A
28 GHz ToA 0.217 0.336 INF
78 GHz ToA 0.197 0.649 INF
Dataset3 Baseline Failed N/A
28 GHz ToA 0.063 0.128 INF
78 GHz ToA 0.063 0.123 INF
Dataset4 Baseline Failed N/A
28 GHz ToA 0.102 0.314 INF
78 GHz ToA 0.102 0.150 INF
Mode Average Baseline Failed N/A
28 GHz ToA 0.118 0.280 INF
78 GHz ToA 0.112 0.320 INF
$\dagger$ Dataset2’s RGB-D Inertial with ToA results were excluded from average calculations due to anomalous degradation.
* The symbol INF indicates cases where ToA enabled successful tracking when the baseline failed.
$\ddagger$ N/A denotes baseline systems without global positioning capability.
V-A Evaluation Methodology: Local vs. Global Alignment
Before delving into the analysis, it is crucial to understand the difference between local and global evaluations and the implications of including baseline global results.
In the local evaluation, the transformation (rotation and translation) between the estimated trajectory and the ground truth is assumed to be unknown and is calculated offline after the SLAM process. This alignment is typically performed using methods like Horn’s absolute orientation algorithm, which finds the best-fit transformation that minimizes the trajectory error. Essentially, the local evaluation focuses on the consistency of the trajectory shape, disregarding the absolute positioning and orientation in the global frame.

### v5

@@ IV-A Experimental Setup and Data Collection
Dataset Frequency (GHz) Base Station Std (cm) Mean (cm) AL01 28 1 $32.92$ $4.20$ 2 $36.05$ $-1.21$ 3 $31.99$ $194$ $41.32$ $8.28$ 78 1 $19.58$ $-2.71$ 2 $19.44$ $1.34$ 3 $19.10$ $0.31$ 4 $14.25$ $-1.72$ AL02 28 1 $36.31$ $-6.88$ 2 $29.60$ $-3.56$ 3 $36.60$ $1.88$ 4 $32.22$ $-2.68$ 78 1 $15.95$ $-2.83$ 2 $17.01$ $-2.29$ 3 $16.08$ $0.56$ 4 $15.64$ $-0.86$ AL03 28 1 $38.44$ $-8.24$ 2 $31.94$ $-4.93$ 3 $33.25$ $7.97$ 4 $30.49$ $-0.50$ 78 1 $15.83$ $0.87$ 2 $18.92$ $-4.97$ 3 $16.38$ $-0.80$ 4 $16.63$ $-0.59$ AL04 28 1 $40.14$ $-14.56$ 2 $32.19$ $-6.68$ 3 $31.17$ $494$ $27.64$ $-4.97$ 78 1 $15.03$ $0.87$ 2 $17.61$ $-8.71$ 3 $14.69$ $-1.49$ 4 $15.94$ $-1.06$ AL05 28 1 $38.23$ $-19.41$ 2 $29.54$ $-2.18$ 3 $31.23$ $-3.09$ 4 $40.62$ $-8.31$ 78 1 $14.40$ $3.12$ 2 $16.80$ $-4.67$ 3 $15.56$ $0.57$ 4 $18.76$ $-1.08$
Table I shows the statistics of the simulated ToA measurements. The 78 GHz configuration demonstrates superior precision with standard deviations of 14.25-19.58 cm compared to 27.64-41.32 cm for 28 GHz. Mean errors at 78 GHz typically remain within ±5 cm, while 28 GHz exhibits larger biases up to 19.41 cm. This performance advantage can be attributed to the larger bandwidth and higher subcarrier spacing available at 78 GHz.
IV-C Evaluation Metrics
We evaluated our approach on both the Aerolab and EuRoC MAV datasets across multiple SLAM configurations (RGB-D, RGB-D-Inertial, Monocular, Monocular-Inertial, and Stereo) using the Absolute Trajectory Error (ATE) metric. For each configuration, we report:
  • Local error (meters): Trajectory accuracy after SE3 post-alignment (rotation and translation).
  • Global error (meters): Absolute positioning accuracy without post-alignment.
For monocular configurations, which suffer from scale ambiguity, we additionally report:
  • Local error: Trajectory accuracy after Sim3 alignment (rotation, translation, and scale).
  • Unscaled local error: Trajectory accuracy with only SE3 alignment (no scale correction).
  • Scale error (%): Percentage deviation from true scale.
In local evaluation, the transformation between estimated trajectory and ground truth is calculated offline after SLAM completion, focusing on trajectory shape consistency. Global evaluation operates without post-processing alignment, requiring the system to maintain accurate absolute positioning throughout operation. To ensure statistical reliability, we conducted each experiment five times and reported the average ATE values across these runs.
@@ V Results
V-A Global SLAM Performance with 5G ToA Integration
This section presents the experimental results evaluating global SLAM performance with 5G ToA integration at two frequencies: 28 GHz and 78 GHz. We evaluate the system on two distinct datasets:
1. The Aerolab dataset: Performance metrics for RGB-D, RGB-D Inertial, and Monocular SLAM configurations (detailed in Table II).
2. The EuRoC MAV dataset: Stereo and Monocular SLAM performance across five sequences (V101-V103, V201-V202) (detailed in Table III).
The tables include baseline measurements (without ToA integration) and results with ToA integration at both frequency bands, along with mode averages for each SLAM configuration.
TABLE II: Global SLAM Performance with 5G ToA Integration on Aerolab Dataset.
Note: For monocular mode with ToA, Datasets AL03 and AL05 results are calculated up to 75s and 65s, as tracking was lost beyond these points.
(a) RGB-D SLAM (SE3 Post-Alignment)
Dataset Configuration Local (m) Global (m) Local Imp.
AL01 Baseline 0.064 N/A
@@ V-A Global SLAM Performance with 5G ToA Integration
(b) RGB-D Inertial SLAM (SE3 Post-Alignment)
Dataset Configuration Local (m) Global (m) Local Imp.
AL01 Baseline 0.089 N/A
@@
28 GHz ToA 0.94 0.280 0.42%
78 GHz ToA 0.90 0.262 5.97%
(c) Monocular SLAM (Both SE3 and Sim3 Post-Alignment for ToA)
Dataset Configuration Local (m) Global (m) Local (Unscaled) Scale Local Imp.
AL01 Baseline Failed N/A N/A N/A
@@ 78 GHz ToA Failed Failed
AL02 Baseline 0.055 N/A N/A N/A
28 GHz ToA 0.055 0.380 0.059 0.93% 0.6%
78 GHz ToA 0.057 0.121 0.058 0.37% -3.3%
AL03 Baseline 0.058 N/A N/A N/A
28 GHz ToA 0.084 0.186 0.089 1.58% -44.0%
78 GHz ToA 0.074 0.152 0.075 0.62% -26.6%
AL04 Baseline 0.074 N/A N/A N/A
28 GHz ToA 0.077 0.162 0.084 1.97% -3.8%
78 GHz ToA 0.056 0.127 0.058 0.76% 23.8%
AL05 Baseline 0.062 N/A N/A N/A
28 GHz ToA 0.075 0.260 0.084 2.44% -20.2%
78 GHz ToA 0.054 0.131 0.061 1.90% 13.8%
Mode Average Baseline 0.064 N/A N/A N/A
28 GHz ToA 0.069 0.267 0.076 0.30% -7.8%
78 GHz ToA 0.056 0.126 0.059 0.15% 11.4%
$\dagger$ AL03’s RGB-D Inertial and monocular ToA results were excluded from average due to anomalous degradation.
$\ddagger$ N/A denotes baseline systems without global positioning and scaling capability.
Monocular-inertial configurations failed across all datasets due to initialization issues, preventing successful tracking.
For RGB-D and RGB-D Inertial modes, accuracy is evaluated using SE3 post-alignment.
For Monocular mode, the baseline uses Sim3 alignment, while ToA configurations provide both SE3 and Sim3 alignment results.
TABLE III: Average Errors and Improvements for Different SLAM Modes on EuRoC MAV Dataset.
Note: Results for V203 are excluded due to tracking failures.
(a) Stereo SLAM (SE3 Post-Alignment for ToA)
Sequence Configuration Local (m) Global (m) Local Imp.
V101 Baseline 0.084 N/A
28 GHz ToA 0.084 0.232 0.0%
78 GHz ToA 0.081 0.183 3.6%
V102 Baseline 0.099 N/A
28 GHz ToA 0.113 0.585 -14.1%
78 GHz ToA 0.106 0.139 -7.2%
V103 Baseline 0.161 N/A
28 GHz ToA 0.182 0.359 -13.0%
78 GHz ToA 0.126 0.190 21.7%
V201 Baseline 0.113 N/A
28 GHz ToA 0.094 0.335 17.2%
78 GHz ToA 0.086 0.177 23.8%
V202 Baseline 0.190 N/A
28 GHz ToA 0.196 0.274 -3.0%
78 GHz ToA 0.192 0.265 -1.0%
Average Baseline 0.129 N/A
28 GHz ToA 0.134 0.357 -2.6%
78 GHz ToA 0.118 0.191 8.2%
(b) Monocular SLAM (SE3 and Sim3 Post-Alignment for ToA)
Sequence Configuration Local (m) Global (m) Local (Unscaled) (m) Scale Local Imp.
V101 Baseline 0.095 N/A N/A N/A
28 GHz ToA 0.094 0.247 0.099 1.55% 1.1%
78 GHz ToA 0.094 0.149 0.094 0.05% 1.1%
V102 Baseline 0.100 N/A N/A N/A
28 GHz ToA 0.100 0.438 0.102 0.40% 0.0%
78 GHz ToA 0.098 0.128 0.098 0.59% 2.0%
V103 Baseline 0.116 N/A N/A N/A
28 GHz ToA 0.103 0.224 0.111 2.24% 11.2%
78 GHz ToA 0.102 0.415 0.102 0.01% 12.1%
V201 Baseline 0.086 N/A N/A N/A
28 GHz ToA 0.085 0.669 0.125 3.75% 1.2%
78 GHz ToA 0.085 0.247 0.085 0.17% 1.2%
V202 Baseline 0.176 N/A N/A N/A
28 GHz ToA 0.178 0.579 0.179 0.74% -1.1%
78 GHz ToA 0.178 0.248 0.181 1.36% -1.1%
Average Baseline 0.115 N/A N/A N/A
28 GHz ToA 0.112 0.431 0.123 1.74% 2.5%
78 GHz ToA 0.111 0.237 0.112 0.20% 3.1%
Before delving into the analysis, it is crucial to understand the difference between local and global evaluations and the implications of including baseline global results. In the local evaluation, the transformation (rotation and translation) between the estimated trajectory and the ground truth is assumed to be unknown and is calculated offline after the SLAM process. This alignment is typically performed using methods like Horn’s absolute orientation algorithm, which finds the best-fit transformation that minimizes the trajectory error. Essentially, the local evaluation focuses on the consistency of the trajectory shape, disregarding the absolute positioning and orientation in the global frame.